2. Consider the information in Table 1 Table 1 Correlation with market portfolio 0.20 0.80 1.00 0.00 Standard deviation Return Beta Stock 1 Stock 2 Market portfolio 6% 12% 8% 0% 16% 2% Risk-free asset 0 (a) Consider Table 1. Calculate betas for stock 1 and stock 2. (b) Consider Table 1. Compute the equilibrium expected return according to the CAPM for stocks 1 and 2. (c) Consider Table 1 and the equilibrium expected returns computed in part (b) of this question. Suppose the correlation between the returns on stocks 1 and 2 is 0.16. Compute the expected return and standard deviation of return of an equally-weighted portfolio of stocks 1 and 2. (d) Consider Table 1. Graphically illustrate various combinations of portfolio risk and return that can be generated by investing in stocks 1 and 2. Consider Table I. Suppose that a risk-free asset is available and offers a certain return of 2%. Solve for the composition, the expected return and standard deviation of the tangency portfolio. Sketch the set of portfolios consisting of the risk-free asset and the tangency portfolio (e)

I would like part d and e answered please
Answer #1

a.

Calculating Beta,

Beta = (Correlation*Standard deviation of stock)/Standard Deviation of market portfolio

Beta for Stock 1 = (0.2*0.06)/0.08

Beta for Stock 1 = 0.15

Beta for Stock 2 = (0.8*0.12)/0.08

Beta for Stock 2 = 1.2

b.

Calculating equilibrium return using CAPM model,

r = Rf + Beta(Rm - Rf )

Equilibrium Return for Stock 1 = 0.02 + 0.15(0.16 - 0.02)

Equilibrium Return for Stock 1 = 4.1%

Equilibrium Return for Stock 2 = 0.02 + 1.2(0.16 - 0.02)

Equilibrium Return for Stock 2 = 18.8%

c.

Formula for Expected Return for equally weighted portfolio,

Expected Return = 0.5(R1) + 0.5(R2)

Expected Return = 0.5*4.1 + 0.5*18.8

Expected Return = 2.05 + 9.4

Expected Return = 11.45

Covariance = Correlation*(Standard Deviation of 1)*(Standard Deviation of 2)

Covariance = 0.16*6*12

Covariance = 11.52

Portfolio Standard Deviation= (w2A2(RA) + w2B2(RB) + 2*(wA)*(wB)*Cov(RA, RB) )0.5

Formula for Standard Deviation of Portfolio

Here, portfolio is equally weighted so w = 0.5 and putting values of standard deviation from table

Portfolio Standard Deviation = 7.12%